# Principal Component Copulas for Capital Modeling and Systemic Risk

### v2

@@ 5 Conclusion
We have applied our techniques to the international stock market in order to study global diversification and systemic risk. We have shown that PCCs provide clear economic interpretations for their copula generators in terms of collective movements in the global market. Recent factor copula studies have either assumed that all factors have a common stochastic mixing variable, or that all factors are independent. We have proposed a PCC that is based on two stochastic mixing variables, namely one for the parallel direction and one for the orthogonal directions. This leads to enhanced performance on information criteria (AIC and BIC) and measures of systemic risk due to the ability to distinguish between parallel market movements, which enhance systemic risk, and orthogonal movements, which reduce systemic risk. We have expanded existing studies of systemic risk by introducing the notion of the market distress cube, in order to study systemic risk in three directions, namely more extreme crashes (quantile $q$), a larger market (dimension $d$), and increasing market share (fraction $k/d$). Finally, we have developed a binomial test to assess the ability of copulas to capture high-dimensional dependence. In contrast to the PCC we find that commonly used copulas in finance fail this test for systemic risk, even when they perform well on pairwise dependence measures.
@@
Acknowledgements
We are grateful to our colleagues Peter Verhoog and Pascal Lieshout, for valuable discussions on the Principal Component Copula model. The views expressed in this paper are the personal views of the authors and do not necessarily reflect the views or policies of their current or past employers. The authors have no competing interests.

### v3

@@ 5 Conclusion
We have applied our techniques to the international stock market in order to study global diversification and systemic risk. We have shown that PCCs provide clear economic interpretations for their copula generators in terms of collective movements in the global market. Recent factor copula studies have either assumed that all factors have a common stochastic mixing variable, or that all factors are independent. We have proposed a PCC that is based on two stochastic mixing variables, namely one for the parallel direction and one for the orthogonal directions. This leads to enhanced performance on information criteria (AIC and BIC) and measures of systemic risk. We have expanded existing studies of systemic risk by introducing the notion of the market distress cube, in order to study systemic risk in three directions, namely more extreme crashes (quantile $q$), a larger market (dimension $d$), and increasing market share (fraction $k/d$). Finally, we have developed a binomial test to assess the ability of copulas to capture high-dimensional dependence. We find that PCCs pass such binomial tests, while commonly used copulas in finance fail to capture systemic risk, even when they perform well on pairwise dependence measures.